# Robust Regret Optimal Control

### v4

@@ II-A Notation
Let $\mathbb{R}^{n}$ and $\mathbb{R}^{n\times m}$ denote the sets of real $n\times 1$ vectors and $n\times m$ matrices, respectively. Similarly, $\mathbb{C}^{n}$ and $\mathbb{C}^{n\times m}$ denote the sets of complex vectors and matrices of given dimensions. The superscripts $\top$ and *** denote transpose and complex conjugate (Hermitian) transpose of a matrix. Moreover, if $M\in\mathbb{C}^{n\times n}$ then $M^{-\top}$ denotes $(M^{\top})^{-1}=(M^{-1})^{\top}$. The Euclidean norm (2-norm) for a vector $v\in\mathbb{C}^{n}$ is defined to be $\|v\|_{2}:=\sqrt{v^{*}v}=\sqrt{\sum_{i=1}^{n}|v_{i}|^{2}}$. The induced 2-norm for a matrix $M\in\mathbb{C}^{n\times m}$ is defined to be:
@@
The matrix $A\in\mathbb{R}^{n\times n}$ is said to be Schur stable if the spectral radius is $<1$. If $A$ is Schur stable then $G$ is a causal, stable system. Hence $G$ maps an input $d\in\ell_{2}$ to output $e\in\ell_{2}$ starting from the initial condition $x_{-\infty}=0$. The induced $\ell_{2}$-norm for a stable system $G$ is defined to be:
@@ II-C Optimal Non-Causal Controller
Assume $(A,B_{u},C_{e},D_{eu})$ satisfy conditions $(i)-(iv)$ in Lemma 1. Let $X\succeq 0$ be the unique stabilizing solution of the DARE with corresponding gain $K_{x}$.
@@ II-D Output Feedback Control Design
$\hat{A}:=\begin{bmatrix}\hat{A}_{11}&-B_{u}K_{v}\hat{A}_{11}^{-\top}\\ 0&\hat{A}_{11}^{-\top}\end{bmatrix},$ $\hat{B}:=\begin{bmatrix}B_{d}\\ -XB_{d}\end{bmatrix},$
@@
Assume that $(A,B_{u},C_{e},D_{eu})$ satisfy conditions $(i)-(iv)$ in Lemma 1 so that the DARE (22) has a stabilizing solution $X\succeq 0$. This stabilizing solution is used to define the dynamics $\hat{P}$ from $d_{t}$ to $\hat{e}_{t}$ in (29).
@@ III-A Problem Formulation
Let $F_{U}(P,\Delta)$ denote the system obtained by closing $\Delta\in\mathbf{\Delta}$ around the top channels of $P$.§§§The system $F_{L}(F_{U}(P,\Delta),K)$ corresponds to closing a controller $K$ around the bottom channels of $F_{U}(P,\Delta)$. Thus $F_{L}(F_{U}(P,\Delta),K)$ is equivalent to the shorter notation $CL(P,K,\Delta)$. The “nominal” (or best-guess) model is obtained with $\Delta=0$. The results in Section II can be used to construct a controller that achieves $(\gamma_{d},\gamma_{J})$-regret on the nominal model $F_{U}(P,0)$. In this section we focus on designing a single controller that achieves a certain level of regret for all possible models in the set :=$\mathcal{M}:=\{F_{U}(P,\Delta)\,:\,\Delta\in\mathbf{\Delta}\}$.
@@ III-B Output Feedback Control Design
$\bar{\sigma}\left(\left[\begin{smallmatrix}D(\theta)\cdot I_{n_{v}}&0\\ 0&I_{n_{e}}\end{smallmatrix}\right]M(e^{j\theta})\left[\begin{smallmatrix}D(\theta)^{-1}\cdot I_{n_{w}}&0\\ 0&I_{n_{d}}\end{smallmatrix}\right]\right)<1$ (62)
@@ IV Examples
This section presents two examples to illustrate the proposed synthesis method for robust, regret optimal control. The code to reproduce the results is available at:
@@ IV-A Simple SISO Design
where $W_{unc}(s)=\frac{3s+4.62}{s+23.1}$ and $\Delta(s)$ is a stable, LTI system satisfying $\|\Delta\|_{\infty}\leq 1$. The weight $W_{unc}(s)$ has DC gain of 0.2, $|W_{unc}(j\infty)|=3$, and a zero near $-1.5$rad/sec. This represents a relatively small uncertainty (20%) uncertainty at low frequencies with increasing uncertainty above $1.5$rad/sec. The continuous-time plant and actuator models are discretized using zero-order hold with a sample time $T_{s}=0.001$ sec.$A_{0}(s)$ and $W_{unc}(s)$ are discretized with sample time $T_{s}$. The normalized uncertainty is replaced by a discrete-time, stable LTI system $\Delta(z)$ with sample time $T_{s}$ satisfying $\|\Delta\|_{\infty}\leq 1$.
@@
[…]
@@
Next we consider the nominal regret, as defined in Section II, with $\Delta=0$. The nominal regret includes three special optimal controllers (values within numerical tolerances): (a) $H_{\infty}$: $(\gamma_{d},\gamma_{J})=(1.82,0)$, (b) competitive ratio: $(\gamma_{d},\gamma_{J})=(0,3.29)$, and (c) additive regret: $(\gamma_{d},\gamma_{J})=(1.63,1)$. Figure 7 shows these three points along with the Pareto optimal curve for feasible values $(\gamma_{d},\gamma_{J})$ of nominal regret (blue dashed curve). Equation 21 in Section II-B characterized the Pareto front using a parameterized optimization. Here we use a different numerical implementation and generate the curve by: (i) selecting 20 values of $\gamma_{d}$ in the range of [0.001, $[0.001,\,0.999]\times\gamma_{\infty}$ where $\gamma_{\infty}=1.82$, and (ii) bisecting to find the minimal value of $\gamma_{J}$ for each value of $\gamma_{d}$ in the grid. Step 2 was performed using the method described in Section II. The bisection was stopped when the interval of lower/upper bounds on $\gamma_{J}$, denoted $[\underline{\gamma}_{J},\bar{\gamma}_{J}]$, satisfied $\bar{\gamma}_{J}-\underline{\gamma}_{J}\leq 10^{-4}+10^{-4}\bar{\gamma}_{J}$. The twenty points were computed in $\approx 6.4$sec on a standard laptop, i.e. each point on the nominal regret curve took $\approx 0.32$sec to compute.
@@ IV-B Boeing 747
Figure 10 shows these three points along with the Pareto optimal curve for feasible values $(\gamma_{d},\gamma_{J})$ of nominal regret (blue dashed curve). The curve was generated by: (i) selecting 20 values of $\gamma_{d}$ in the range of [0.001, $[0.001,\,0.999]\gamma_{\infty}$ where $\gamma_{\infty}=28.47$, and (ii) bisecting to find the minimal value of $\gamma_{J}$ for each value of $\gamma_{d}$ in the grid. Step 2 was performed using the method described in Section II. The bisection was stopped when the interval of lower/upper bounds on $\gamma_{J}$, denoted $[\underline{\gamma}_{J},\bar{\gamma}_{J}]$, satisfied $\bar{\gamma}_{J}-\underline{\gamma}_{J}\leq 10^{-2}+10^{-3}\bar{\gamma}_{J}$. The twenty points were computed in $\approx 4.5$sec on a standard laptop, i.e. each point on the nominal regret curve took $\approx 0.45$sec to compute.
@@ IV-C Quarter Car Example
[…]
@@
where $m_{b}=300$kg, $m_{w}=60$kg, $b_{s}=1000$N/m/s, $k_{s}=16000$N/m, and $k_{t}=1.9\times 10^{5}$/m. The model has lightly complex poles at $-1.43\pm 6.91j$ and $-8.57\pm 57.6j$. The goal is to design the active suspension controller to balance driver comfort (small $a_{b}$) and road handling (small $s_{d}$).
@@
where $W_{unc}(s)=\frac{3s+18.5}{s+46.3}$ and $\Delta(s)$ is a stable, LTI system satisfying $\|\Delta\|_{\infty}\leq 1$. The weight $W_{unc}(s)$ has DC gain of 0.4, $|W_{unc}(j\infty)|=3$, and a zero near $-6.2$rad/sec. This represents a relatively small uncertainty (40%) uncertainty at low frequencies with increasing uncertainty above $6.2$rad/sec. The continuous-time quarter car and hydraulic actuator models are discretized using zero-order hold and sample time $T_{s}=0.002$ sec.$A_{0}(s)$ and $W_{unc}(s)$ are discretized with sample time $T_{s}$. The normalized uncertainty is replaced by a discrete-time, stable LTI system $\Delta(z)$ with sample time $T_{s}$ satisfying $\|\Delta\|_{\infty}\leq 1$.
@@
Figure 13 shows a closed-loop, time-domain response with the nominal additive regret controller. The road disturbance is $r(t)=0.025\cdot(1-\cos(8\pi t))$ for $0\leq t\leq 0.2$ and $r(t)=0$ for $t>0.2$. The simulations are performed with the nominal actuator model and 50 samples of the uncertain actuator model. The nominal responses are good for both the body acceleration and suspension displacement. The displacement remains within the allowed ±$\pm 0.05$m of travel. However, the simulations with the uncertain actuator show a wide variability including lightly damped, oscillatory responses.

### v5

@@ II-A Notation
Let $\mathbb{R}^{n}$ and $\mathbb{R}^{n\times m}$ denote the sets of real $n\times 1$ vectors and $n\times m$ matrices, respectively. Similarly, $\mathbb{C}^{n}$ and $\mathbb{C}^{n\times m}$ denote the sets of complex vectors and matrices of given dimensions. The superscripts $\top$ and * denote transpose and complex conjugate (Hermitian) transpose of a matrix. Moreover, if $M\in\mathbb{C}^{n\times n}$ then $M^{-\top}$ denotes $(M^{\top})^{-1}=(M^{-1})^{\top}$. The Euclidean norm (2-norm) for a vector $v\in\mathbb{C}^{n}$ is defined to be $\|v\|_{2}:=\sqrt{v^{*}v}=\sqrt{\sum_{i=1}^{n}|v_{i}|^{2}}$. The induced 2-norm for a matrix $M\in\mathbb{C}^{n\times m}$ is defined to be:
@@
The matrix $A\in\mathbb{R}^{n\times n}$ is said to be Schur stable if the spectral radius is <1<1. If $A$ is Schur stable then $G$ is a causal, stable system. Hence $G$ maps an input $d\in\ell_{2}$ to output $e\in\ell_{2}$ starting from the initial condition $x_{-\infty}=0$. The induced $\ell_{2}$-norm for a stable system $G$ is defined to be:
@@ II-C Optimal Non-Causal Controller
Assume $(A,B_{u},C_{e},D_{eu})$ satisfy conditions (i)(iv)(i)-(iv) in Lemma 1. Let $X\succeq 0$ be the unique stabilizing solution of the DARE with corresponding gain $K_{x}$.
@@ II-D Output Feedback Control Design
$\hat{A}:=\begin{bmatrix}\hat{A}_{11}&-B_{u}K_{v}\hat{A}_{11}^{-\top}\\ 0&\hat{A}_{11}^{-\top}\end{bmatrix}$, $\hskip-7.22743pt\hat{B}:=\begin{bmatrix}B_{d}\\ -XB_{d}\end{bmatrix}$,
@@
Assume that $(A,B_{u},C_{e},D_{eu})$ satisfy conditions (i)(iv)(i)-(iv) in Lemma 1 so that the DARE (7) has a stabilizing solution $X\succeq 0$. This stabilizing solution is used to define the dynamics $\hat{P}$ from $d_{t}$ to $\hat{e}_{t}$ in (14).
@@ III-A Problem Formulation
Let $F_{U}(P,\Delta)$ denote the system obtained by closing $\Delta\in\mathbf{\Delta}$ around the top channels of PP.§§§The system $F_{L}(F_{U}(P,\Delta),K)$ corresponds to closing a controller $K$ around the bottom channels of $F_{U}(P,\Delta)$. Thus $F_{L}(F_{U}(P,\Delta),K)$ is equivalent to the shorter notation $CL(P,K,\Delta)$. The “nominal” (or best-guess) model is obtained with $\Delta=0$. The results in Section II can be used to construct a controller that achieves $(\gamma_{d},\gamma_{J})$-regret on the nominal model $F_{U}(P,0)$. In this section we focus on designing a single controller that achieves a certain level of regret for all possible models in the set $\mathcal{M}:=\{F_{U}(P,\Delta)\,:\,\Delta\in\mathbf{\Delta}\}$.
@@ III-B Output Feedback Control Design
$\bar{\sigma}\left(\left[\begin{smallmatrix}D(\theta)\cdot I_{n_{v}}&0\\ 0&I_{n_{e}}\end{smallmatrix}\right]M(e^{j\theta})\left[\begin{smallmatrix}D(\theta)^{-1}\cdot I_{n_{w}}&0\\ 0&I_{n_{d}}\end{smallmatrix}\right]\right)<1$
@@ IV-A Simple SISO Design
[…]
@@
$J(K^{o},d)$ $=\frac{1}{2\pi}\int_{-\infty}^{+\infty}\hat{e}(j\omega)^{*}\,\hat{e}(j\omega)\,d\omega=\frac{1}{2\pi}\int_{-\infty}^{+\infty}\left\|T^{o}(j\omega)\right\|_{2}^{2}\,\,|\hat{d}(j\omega)|^{2}\,d\omega.$ (29)
@@
Next we consider the nominal regret, as defined in Section II, with $\Delta=0$. The nominal regret includes three special optimal controllers (values within numerical tolerances): (a) $H_{\infty}$: $(\gamma_{d},\gamma_{J})=(1.82,0)$, (b) competitive ratio: $(\gamma_{d},\gamma_{J})=(0,3.29)$, and (c) additive regret: $(\gamma_{d},\gamma_{J})=(1.63,1)$. Figure 7 shows these three points along with the Pareto optimal curve for feasible values $(\gamma_{d},\gamma_{J})$ of nominal regret (blue dashed curve). Equation 6 in Section II-B characterized the Pareto front using a parameterized optimization. Here we use a different numerical implementation and generate the curve by: (i) selecting 20 values of $\gamma_{d}$ in the range of $0.001,\,0.999]\times\gamma_{\infty}$ where $\gamma_{\infty}=1.82$, and (ii) bisecting to find the minimal value of $\gamma_{J}$ for each value of $\gamma_{d}$ in the grid. Step 2 was performed using the method described in Section II. The bisection was stopped when the interval of lower/upper bounds on $\gamma_{J}$, denoted $\underline{\gamma}_{J},\bar{\gamma}_{J}]$, satisfied $\bar{\gamma}_{J}-\underline{\gamma}_{J}\leq 10^{-4}+10^{-4}\bar{\gamma}_{J}$. The twenty points were computed in $\approx 6.4$sec on a standard laptop, i.e. each point on the nominal regret curve took $\approx 0.32$sec to compute.
@@ IV-B Boeing 747
Figure 10 shows these three points along with the Pareto optimal curve for feasible values $(\gamma_{d},\gamma_{J})$ of nominal regret (blue dashed curve). The curve was generated by: (i) selecting 20 values of $\gamma_{d}$ in the range of $0.001,\,0.999]\gamma_{\infty}$ where $\gamma_{\infty}=28.47$, and (ii) bisecting to find the minimal value of $\gamma_{J}$ for each value of $\gamma_{d}$ in the grid. Step 2 was performed using the method described in Section II. The bisection was stopped when the interval of lower/upper bounds on $\gamma_{J}$, denoted $\underline{\gamma}_{J},\bar{\gamma}_{J}]$, satisfied $\bar{\gamma}_{J}-\underline{\gamma}_{J}\leq 10^{-2}+10^{-3}\bar{\gamma}_{J}$. The twenty points were computed in $\approx 4.5$sec on a standard laptop, i.e. each point on the nominal regret curve took $\approx 0.45$sec to compute.
@@ IV-C Quarter Car Example
where $r(t)\in\mathbb{R}$ is the road disturbance (m), $f_{s}(t)\in\mathbb{R}$ is the active suspension force (kN), $a_{b}(t)\in\mathbb{R}$ is the car body acceleration (m/s2), $s_{d}(t)\in\mathbb{R}$ is the suspension displacement (m), and $x(t)\in\mathbb{R}^{4}$ is the state. The state-space data is:
@@
where $m_{b}=300$kg, $m_{w}=60$kg, $b_{s}=1000$N/m/s, $k_{s}=16000$N/m, and $k_{t}=1.9\times 10^{5}$N/m. The model has lightly complex poles at $1.43\pm 6.91j$ and $8.57\pm 57.6j$. The goal is to design the active suspension controller to balance driver comfort (small $a_{b}$) and road handling (small $s_{d}$).
@@
[…]
@@
Figure 13 shows a closed-loop, time-domain response with the nominal additive regret controller. The road disturbance is $r(t)=0.025\cdot(1-\cos(8\pi t))$ for $0\leq t\leq 0.2$ and $r(t)=0$ for $t>0.2$. The simulations are performed with the nominal actuator model and 50 samples of the uncertain actuator model. The nominal responses are good for both the body acceleration and suspension displacement. The displacement remains within the allowed $\pm 0.05$m of travel. However, the simulations with the uncertain actuator show a wide variability including lightly damped, oscillatory responses.